OpenStudy (anonymous):

I need help guys 2. Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 3.If g(x) = x2 + 3, find g(4). 4.If g(x) = 2(x - 4), find the value of x if g(x) = 20. 5. For the function f(x) = x + 7, what is the ordered pair for the point on the graph when x = 2b?

OpenStudy (anonymous):

@bibby @Skyz @Ashleyisakitty @kewlgeek555 @Luigi0210 @♪Chibiterasu @wolfe8 @shamil98 @austinL

OpenStudy (bibby):

The first bit is recursive. the sequence starts off 2,3... the third number is the sum of the previous 2. so here, 2,3,(2+3),(2+3+3) etc. 2,3,5,8,13

OpenStudy (the_fizicx99):

bibby's got this. Byee

OpenStudy (bibby):

plug in 4 into g(x) and solve\[4^2+3\] and the other\[ g(20)=2(20-4)\] I love you fizic

OpenStudy (the_fizicx99):

Luh you.

OpenStudy (anonymous):

Omg bibs you are a life saver

OpenStudy (bibby):

@nincompoop get on number 5 pls as I do not understand it

OpenStudy (bibby):

unless b = 7 cause of y = mx + b in which case x = 14 and yeah

OpenStudy (the_fizicx99):

y = mx + b ?? >.>

OpenStudy (nincompoop):

yep

OpenStudy (the_fizicx99):

Wouldn't 7 be your B in y = mx + b?

OpenStudy (nincompoop):

it is an equation of a line y = mx+b so substitute x = 2b = 2(7) = 14

OpenStudy (the_fizicx99):

Yea, its 14

OpenStudy (anonymous):

These are the choices for 5 (2b, 2b + 7) (2b, x + 7) (x, x + 7) (x, 2b + 7)

OpenStudy (nincompoop):

b is your y-intercept (0,y)

OpenStudy (the_fizicx99):

B ?

OpenStudy (anonymous):

woohoo 90% thx guys

OpenStudy (the_fizicx99):

:o

OpenStudy (nincompoop):

what? you didn't even answer number 5 lol

OpenStudy (anonymous):

I figured it out ^-^ thx love u guys

OpenStudy (nincompoop):

okay